Session: Nov 1997 (session 4)
Wavelength: 1.3 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina ( at Wed Nov 19 01:37:53 MET 1997) Minor failures
it seem all OK.. Weather fog. bye

Effelsberg (dgraham at Fri Nov 21 09:44:45 MET 1997) Minor failures
Damp weather. Tsys values scatter.

Urumqi (Li Suqin at Fri Nov 21 12:00:27 MET 1997) Minor failures
It is cloudy. H-maser to GPS offset is 3.12us. Format to GPS offset is 3.01us. From 06:16:00 to 16:15:35 because the receiver trouble no data be recorded. After 16:15:35 the receiver was not stable.

Jodrell (Alastair Gunn at Sat Nov 22 19:41:53 MET 1997) Minor failures
Due to serious power failure we did not begin observing this schedule until the beginning of tape 2. First tape does not contain any data. Some attenuation changes, otherwise no other known problems.

Metsahovi (Kaj Wiik at Mon Nov 24 13:35:10 MET 1997) Minor failures
WX mostly clear. Problem with unloading the tape, missed some of the first scan but second scan should be OK (forwarded the tape to correct position). Still spurious ERROR messages due to too long comment line in the procedure file (my fault...).

Noto (C.Trigilio at Tue Nov 25 15:41:29 MET 1997) Success
No problem reported. Sky cover 40% at 7UT; 100% at 14 UT, clear at end.

Calibration (Fredrik T Rantakyro at Tue Dec 9 15:05:35 MET 1997) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the NOV97 directory under the name em22d.antab Read the comments in the file to get further information.